OpenStudy (anonymous):

Find an equation for a line Perpendicular to y=-1/4x + 2 that contains (15,-8) -8 = 4/1(15)+b -8=60+b -60 -60 __________ -68 = b So for my final answer ... when we put it in y = mx+b form.. would it be y = -1/4x - 68 OR y = 4x - 68 ??

OpenStudy (anonymous):

Perpendicular line has slope of 4. So, the line is: y = 4x + b Passing through (15, -8). So, b = -8-60 = -68. So, line is: y = 4x-68.

OpenStudy (anonymous):

oooh ok so the slope never remains the same after u change

OpenStudy (anonymous):

Parallel lines have same slope. So, m1 = m2.

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

Sorry, perpendicular lines have product of slopes as -1. m1*m2 = -1